# InfAlign: Inference-aware language model alignment

### v2

@@ 3 Reinforcement learning with reward transformation
In this section, we propose a general framework for solving the language model alignment problem defined in Definition 5. Our approach is based on designing a new reward function $\mathcal{R}$ based on the reward model $r$, the inference-time procedure $T$, and the base policy ${\pi_{\rm ref}}$, such that solving the KL-regularized RL problem (Definition 1) with the transformed reward $\mathcal{R}$ leads to an almost optimal solution to Eq. 5. More precisely, the aligned policy is the solution to the following optimization problem:
$\max_{\pi}\mathbb{E}_{{\bm{x}}\sim\mu,{\bm{y}}\sim\pi(\cdot\mid{\bm{x}})}\{\mathcal{R}_{r,{\pi_{\rm ref}},T}({\bm{x}},{\bm{y}})-\beta D_{\text{KL}}(\pi(\cdot\mid{\bm{x}})\|{\pi_{\rm ref}}(\cdot\mid{\bm{x}}))\},$ (6)
where $\mathcal{R}_{r,{\pi_{\rm ref}},T}({\bm{x}},{\bm{y}})$ is a transformed reward function. At first, it might not be immediately clear how Eq. 6 might help solve the problem in Eq. 5, however, we will show that for any inference-time procedure $T$, there exists a transformed reward $\mathcal{R}$, which solves Eq. 5.
@@
For any base policy ${\pi_{\rm ref}}$, reward model $r$, inference-time procedure $T$, and $\beta>0$, there exists a reward function $\mathcal{R}_{r,{\pi_{\rm ref}},T}$ such that the solution to Eq. 6 solves the optimization problem in Eq. 5 (Definition 5).
In general, such optimal reward transformation will depend on the base policy ${\pi_{\rm ref}}$, the post-hoc procedure $T$, and the reward function $r$. In the lemma below, we list the property that the reward transformation and the resulting optimal aligned policy must satisfy.
@@
Assuming that $T$ is such that $\partial T(\pi)({\bm{y}}_{1}\mid{\bm{x}})/\partial\pi({\bm{y}}_{2}\mid{\bm{x}})$ exists for all ${\bm{x}},{\bm{y}}_{1},{\bm{y}}_{2}$, then we have the optimal transformed reward $\mathcal{R}$ and the optimal policy $\pi^{*}$ in Eq. 5 must satisfy the following coupled equations $\forall{\bm{x}},{\bm{y}}$
@@
$\mathcal{R}({\bm{x}},{\bm{y}})$ $=\frac{\partial}{\partial\pi({\bm{y}}\mid{\bm{x}})}W^{T}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})$ (8)
$=\sum_{{\bm{z}}}\mathcal{C}_{r,T({\pi_{\rm ref}})}({\bm{x}},{\bm{z}})\frac{\partial}{\partial\pi({\bm{y}}\mid{\bm{x}})}T(\pi)({\bm{z}}\mid{\bm{x}}),$ (9)
where $\mathcal{C}_{r,T({\pi_{\rm ref}})}({\bm{x}},{\bm{z}})$ is the calibrated reward under the inference-time transformed policy.
Missing proofs are presented in Appendix A. Theorem 1 naturally leads to an iterative EM-style algorithm that (I) updates $\pi$ with $\mathcal{R}$ fixed based on Eq. 7 and (II) updates $\mathcal{R}$ with $\pi$ fixed based on Eq. 9 until convergence. However, such algorithm suffers from two drawbacks: first, for general language models, it is inefficient/intractable to evaluate Eq. 9 since it involves evaluating the policy on a large, or even infinite output space; second, it is unclear whether such an algorithm could lead to the optimal solution.
To find more efficient ways to design reward transformations, we examine the case when no inference-time procedure is performed. In this case, $T(\pi)=\pi$ and
$\frac{\partial}{\partial\pi({\bm{y}}\mid{\bm{x}})}T(\pi)({\bm{z}}\mid{\bm{x}})=\mathbbm{1}\left\{{\bm{z}}={\bm{y}}\right\}.$
Eq. 9 will reduce to $\mathcal{R}({\bm{x}},{\bm{y}})=\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})$, the CDF or calibrated reward under ${\pi_{\rm ref}}$.
@@
When no inference-time procedure is performed, i.e. $\forall\pi,T(\pi)=\pi$, the solution to Eq. 6 with $\mathcal{R}({\bm{x}},{\bm{y}})=\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})$ is the solution to Eq. 5.
@@
Note that the above corollary is also observed in Azar et al. (2023); Gui et al. (2024). Hence Theorem 1 can be viewed as a generalization of these results with general inference-time procedures. The observation motivates us to consider a family of reward transformations based on this calibrated reward, described in the next section. As we will see, for the class of calibrated inference-time procedures (Definition 6), different transformations in such family could be efficiently evaluated through a toy language model, which enables the search for good or even optimal transformations.
4 Towards solving InfAlign
We first state a few properties of reward calibration in Section 4.1. Then, in Section 4.2, we demonstrate how this approach enables efficient evaluations of different transformations for calibrated inference-time procedures. We will then use best-of-$N$ and worst-of-$N$ as examples of post-hoc procedures to demonstrate the effectiveness of such an approach in Section 4.3. Missing proofs in the section are presented in Appendix A.
4.1 Reward calibration
Recall the definition of the calibrated reward in Eq. 2. In this section, we assume that $\mathcal{C}_{r,{\pi_{\rm ref}}}$ is already obtained and discuss methods to approximate $\mathcal{C}_{r,{\pi_{\rm ref}}}$ in Section 5. We state a few properties of $\mathcal{C}_{r,{\pi_{\rm ref}}}$ below. The first property states that reward calibration preserves the ordering of the reward. This implies that the win rate evaluated under the calibrated reward stays unchanged.
@@ 4.1 Reward calibration
$r({\bm{x}},{\bm{y}})\geq r({\bm{x}},{\bm{z}})\implies\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\geq\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{z}})$ (10)
Moreover, we have that $\mathcal{C}_{r,{\pi_{\rm ref}}}$ is a canonical transformation of reward, and is invariant under all monotone increasing transformations of the reward function, stated below.
Lemma 3 (The calibrated reward is invariant under monotone increasing transformations of reward).
Let $m:\mathbb{R}\to\mathbb{R}$ be any monotonic increasing function. Then,
$\mathcal{C}_{m(r),{\pi_{\rm ref}}}=\mathcal{C}_{r,{\pi_{\rm ref}}}.$ (11)
In particular, this also immediately implies that $\mathcal{C}_{\mathcal{C}_{r,{\pi_{\rm ref}}},{\pi_{\rm ref}}}=\mathcal{C}_{r,{\pi_{\rm ref}}}.$
This property is useful since as long as the learned reward model $r$ can capture relative human preference between pairs, the calibration of $r$ will be the same. Hence $\mathcal{C}$ is more robust to the learning process of $r$.
@@
The lemma provides us a simple, unified view of the output from a language model through the space of calibrated reward, which is independent from the base policy, and the reward model.
4.2 KL-regularized RL with the calibrated reward
Next we discuss how the calibrated reward can be used in KL-regularized reinforcement learning. As Lemma 4 suggests, after calibration, the reward distribution of the outputs from the base policy is independent from the reward model and the base policy itself. This allows us to design a transformation function $\Phi$, focusing only on the inference-time procedure $T$, to be applied on top of the calibrated reward function, independent of ${\pi_{\rm ref}}$ and $r$.
More precisely, let $\Phi:[0,1]\rightarrow\mathbb{R}$ be a transformation function, we propose the following reward function
$\mathcal{R}_{\Phi}({\bm{x}},{\bm{y}})=\Phi(\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{y}}\mid{\bm{x}})),$ (12)
and we would like the aligned policy to be the solution to the KL-regularized RL problem defined in Definition 1 with reward $\mathcal{R}_{\Phi}({\bm{x}},{\bm{y}})$,
$\pi^{*}_{\mathcal{R}_{\Phi},\beta}(\cdot\mid{\bm{x}})=\arg\max_{\pi}\mathbb{E}_{{\bm{y}}\sim\pi(\cdot\mid{\bm{x}})}\{\mathcal{R}_{\Phi}({\bm{x}},{\bm{y}})\}-\beta D_{\text{KL}}(\pi(\cdot\mid{\bm{x}})\|{\pi_{\rm ref}}(\cdot\mid{\bm{x}})).$ (13)
Inference-aware reward transformation.
For a given inference-time procedure $T$, our goal is to derive or design a suitable transformation $\Phi$, such that the solution leads to a good or even optimal trade-off between the inference-time win rate $W^{T}$ and the KL divergence from the base policy.
Standard win rate (no inference-time procedure).
When no inference-time procedure is employed (i.e., $T$ is the identity mapping), $W^{T}$ reduces to the standard win rate. Setting $\Phi$ to be the identity transformation leads to the optimal win rate vs KL trade-off curve by noting that
$\mathbb{E}_{{\bm{y}}\sim\pi(\cdot\mid{\bm{x}})}\{\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\}=W_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}}).$
In this case, Eq. 13 will be the same as the alignment objective of Azar et al. (2023). Moreover, it can also be shown that when $\Phi(\cdot)=\log(\cdot)$, the solution $\pi^{*}_{\mathcal{R}_{\Phi},\beta}$ recovers the popular best-of-$N$ distillation objective, which has been studied in a recent line of works (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024), and shown to be nearly win rate optimal (Yang et al., 2024; Gui et al., 2024). We also note that while these methods lead to similar optimization objectives, the algorithmic approaches to solve the problems are different. In Section 5, we will discuss a unified algorithm to solve Eq. 13 for general $\Phi$. In Section 6, we compare the results for $\Phi$ set to the identity mapping with the abovementioned baseline approaches.
@@ 4.2 KL-regularized RL with the calibrated reward
An inference-time procedure $T$ is called a calibrated procedure if there exists a mapping function $g_{T}:[0,1]\rightarrow\mathbb{R}$ such that for any $\pi$, $r$, and ${\bm{x}},{\bm{y}}$, we have
$T(\pi)({\bm{y}}\mid{\bm{x}})\propto\pi({\bm{y}}\mid{\bm{x}})\cdot g_{T}(\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}})).$
Our next result shows that for calibrated inference-time procedures, the aligned policy from solving Eq. 13 has a win rate and KL divergence independent of the base policy and reward function.
@@
If $T$ is a calibrated inference-time procedure, for any continuous language model $\pi$, $\beta>0$ and reward transformation function $\Phi$, we have that both $W^{T}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})$ and $D_{\text{KL}}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\|{\pi_{\rm ref}})$ are independent of $r$ and ${\pi_{\rm ref}}$.
The above theorem allows us to evaluate a transformation $\Phi$ by focusing on simple continuous language models that are easy to compute and simulate. In the next section, we will use two popular inference-time procedures, best-of-$N$ and worst-of-$N$, as examples to demonstrate how the theorem enables us to efficiently evaluate the inference-time win rate vs KL divergence tradeoff curve for different $\Phi$ functions, which could be used to find a suitable transformation $\Phi$ in practical scenarios.
4.3 Finding better transformations for BoN and WoN
In this section, we focus on the following two inference-time procedures.
Best-of-$N$ inference-time procedure (BoN). During inference, $N$ i.i.d. responses from a policy $\pi$ are generated, i.e., ${\bm{y}}_{1},\ldots{\bm{y}}_{N}\sim\pi$. The final output is the one with the highest reward, i.e.,
${\bm{y}}_{\texttt{BoN}}=\arg\max_{{\bm{y}}\in\{{\bm{y}}_{1},\ldots{\bm{y}}_{N}\}}r({\bm{x}},{\bm{y}}).$
Worst-of-$N$ inference-time procedure (WoN). Let ${\bm{y}}_{1},\ldots{\bm{y}}_{N}$ be $N$ i.i.d. draws from policy $\pi$. The final output is the one with the lowest reward. i.e.,

### v5

@@ 3 Reinforcement learning with reward transformation
In this section, we propose a general framework for solving InfAlign (Definition 5). Our approach is based on designing a new reward function $\mathcal{R}$ based on the reward model $r$, the inference-time procedure $\mathcal{T}$, and the base policy $\pi_{\rm ref}$}, such that solving the RLHF problem (Definition 1) with the transformed reward $\mathcal{R}$ leads to an optimal solution to InfAlign. More precisely, the aligned policy is the maximizer of the following regularized objective:
$\vskip-10.0pt\mathcal{R}_{r,{\pi_{\rm ref}},\mathcal{T}}({\bm{x}},{\bm{y}})-\beta D_{\text{KL}}(\pi(\cdot\mid{\bm{x}})\|{\pi_{\rm ref}}(\cdot\mid{\bm{x}})),$ (4)
where $\mathcal{R}_{r,{\pi_{\rm ref}},\mathcal{T}}({\bm{x}},{\bm{y}})$ is a transformed reward function. Interestingly, we show that such reward transformation is sufficient to solve InfAlign.
@@
For any base policy $\pi_{\rm ref}$}, reward model $r$, inference-time procedure $\mathcal{T}$, and $\beta>0$, there exists a reward function $\mathcal{R}_{r,{\pi_{\rm ref}},\mathcal{T}}$ such that the maximizer of Eq. 4 solves the optimization problem in Eq. 3 (Definition 5).
In general, such optimal reward transformation will depend on the base policy $\pi_{\rm ref}$}, the inference-time procedure $\mathcal{T}$, and the reward model $r$. In the lemma below, we list the property that the reward transformation and the resulting optimal aligned policy must satisfy.
@@
Assuming that $\mathcal{T}$ is such that $\partial\mathcal{T}_{\pi}({\bm{y}}_{1}\mid{\bm{x}})/\partial\pi({\bm{y}}_{2}\mid{\bm{x}})$ [^4] To make sure that the partial derivative is well-defined, we assume that $\mathcal{T}$ is well-defined for inputs within an infinitesimal expansion of $\Delta_{\bm{\mathcal{Y}}}$. We note that the assumption holds for procedures like BoN and WoN, as shown in Lemma 5. exists for all ${\bm{x}},{\bm{y}}_{1},{\bm{y}}_{2}$, then we have the optimal transformed reward $\mathcal{R}$ and the optimal policy $\pi^{*}$ in Eq. 3 must satisfy the following coupled equations: $\forall{\bm{x}},{\bm{y}}$
@@
$\mathcal{R}({\bm{x}},{\bm{y}})$ $=\frac{\partial}{\partial\pi({\bm{y}}\mid{\bm{x}})}W^{\mathcal{T}}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})|_{\pi=\pi^{*}}$ (6)
$=\sum_{{\bm{z}}}\mathcal{C}_{r,\mathcal{T}_{{\pi_{\rm ref}}}}({\bm{x}},{\bm{z}})\frac{\partial\mathcal{T}_{\pi}({\bm{z}}\mid{\bm{x}})}{\partial\pi({\bm{y}}\mid{\bm{x}})}|_{\pi=\pi^{*}},$ (7)
where the last inequality is due to Eq. 2.
Missing proofs are presented in Appendix B. Theorem 1 naturally leads to an iterative EM-style algorithm that (I) updates $\pi$ with $\mathcal{R}$ fixed based on Eq. 5 and (II) updates $\mathcal{R}$ with $\pi$ fixed based on Eq. 7 until convergence. However, such algorithm suffers from two drawbacks: first, for general language models, it is inefficient/intractable to evaluate Eq. 7 since it involves evaluating the policy on a large, or even infinite output space; second, it is unclear whether such an algorithm could lead to the optimal solution.
To find more efficient ways to design reward transformations, we examine the case when no inference-time procedure is performed. In this case, $\mathcal{T}_{\pi}=\pi$ and
$\frac{\partial}{\partial\pi({\bm{y}}\mid{\bm{x}})}\mathcal{T}_{\pi}({\bm{z}}\mid{\bm{x}})=\mathbbm{1}\left\{{\bm{z}}={\bm{y}}\right\}.$
Eq. 7 will reduce to $\mathcal{R}({\bm{x}},{\bm{y}})=\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})$, the calibrated reward under $\pi_{\rm ref}$}.
@@
When no inference-time procedure is performed, i.e. $\forall\pi,\mathcal{T}_{\pi}=\pi$, the maximizer of Eq. 4 with $\mathcal{R}({\bm{x}},{\bm{y}})=\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})$ is the solution to Eq. 3.
@@
The above corollary is also observed in Azar et al. (2023); Gui et al. (2024). Hence Theorem 1 can be viewed as a generalization of these results with general inference-time procedures. The observation motivates us to consider RLHF with a specific family of reward transformations that involves a reward calibration step, described next.
4 InfAlign-CTRL: Calibrate- and-transform reinforcement learning
In this section, we propose the InfAlign-CTRL method, which is our proposed solver for the InfAlign problem. The method consists of: (1) Calibration: Approximate the calibrated reward $\mathcal{C}_{r,{\pi_{\rm ref}}}$; (2) Transformation: Apply transformation $\Phi$ on top of $\mathcal{C}_{r,{\pi_{\rm ref}}}$ to obtain $\mathcal{R}_{\Phi}=\Phi\circ\mathcal{C}_{r,{\pi_{\rm ref}}}$; (3) Solve RLHF with the transformed reward. We discuss each step in more details below.
Reward calibration.
The goal of this step is to obtain the calibrated reward $\mathcal{C}_{r,{\pi_{\rm ref}}}$ (Eq. 1). We first assume $\mathcal{C}_{r,{\pi_{\rm ref}}}$ could be obtained perfectly and discuss practical ways to approximate it in Section 4.4. In Section 4.1, we discuss various properties of $\mathcal{C}_{r,{\pi_{\rm ref}}}$ In particular, it can be shown that for continuous language models, the distribution of $\mathcal{C}_{r,{\pi_{\rm ref}}}$ is independent from the base policy, and the reward model (Lemma 4), which provides a unified view of the outputs from a language model through the lens of calibrated reward. This allows us to focus the design of the transformation function $\Phi$ on $\mathcal{T}$ in the following step.
As mentioned in Corollary 1, using the calibrated reward directly in the RLHF problem leads to an optimal solution to InfAlign with no inference-time procedure. Note that the resulting solution is the same as the alignment objective of Azar et al. (2023). Moreover, it can also be shown that using $\log\mathcal{C}_{r,{\pi_{\rm ref}}}$ as the reward, the solution to the RLHF problem recovers the popular best-of-$N$ distillation objective, which has been studied in a recent line of works (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024), and shown to be nearly optimal for standard win rate  (Yang et al., 2024; Gui et al., 2024). We note that while these methods lead to similar optimization objectives, InfAlign-CTRL makes this calibration step explicit, resulting in a different algorithmic approach. In Section 5, we compare the results with the abovementioned baseline approaches on standard win rate, and show that it leads to improved win rate. We also empirically show that reward calibration is beneficial to mitigate reward hacking (Appendix D).
Reward transformation.
The goal is to further transform the calibrated reward using a transformation function $\Phi:[0,1]\rightarrow\mathbb{R}$. The function $\Phi$ is chosen based on the inference-time procedure $\mathcal{T}$ so that $\Phi\circ\mathcal{C}_{r,{\pi_{\rm ref}}}$ is a good reward transformation to use for solving InfAlign.
Ideally, we would like the design of $\Phi$ to only depend on the inference-time procedure $\mathcal{T}$ so that it is transferable among different $r$ and $\pi_{\rm ref}$}. A natural question is for what type of $\mathcal{T}$’s this is possible. In Section 4.2, we show that for calibrated inference-time procedures (Definition 6), and continuous language models, it is sufficient for $\Phi$ to depend on $\mathcal{T}$. And in these cases, different $\Phi$’s could be evaluated efficiently with simple language models that can be easily simulated, which enables the search for good or even optimal transformations. We use Best-of-$N$ and Worst-of-$N$ as examples of inference-time procedures to demonstrate the effectiveness of such approach in Section 4.3.
RLHF with the transformed reward.
At this step, we solve the RLHF problem with reward function,
$\mathcal{R}_{\Phi}({\bm{x}},{\bm{y}})=\Phi(\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{y}}\mid{\bm{x}})),$
to obtain the aligned policy $\pi^{*}_{\mathcal{R}_{\Phi},\beta}$. Since we only modify the reward function, the step can be performed with standard RLHF solvers. We present a practical version of the InfAlign-CTRL method in Section 4.4.
4.1 Properties of reward calibration
We present properties of $\mathcal{C}_{r,{\pi_{\rm ref}}}$. The first property states that reward calibration preserves the ordering of the reward.
@@ 4.1 Properties of reward calibration
$r({\bm{x}},{\bm{y}})\geq r({\bm{x}},{\bm{z}})\implies\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\geq\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{z}}).$
Moreover, $\mathcal{C}_{r,{\pi_{\rm ref}}}$ is invariant under all monotone increasing transformations of the reward function, stated below.
Lemma 3 (Calibration is invariant under monotone increasing transformations).
Let $m:\mathbb{R}\to\mathbb{R}$ be any strictly monotone increasing function. Then $\mathcal{C}_{m(r),{\pi_{\rm ref}}}=\mathcal{C}_{r,{\pi_{\rm ref}}}.$
This property is useful since as long as the learned reward model $r$ can capture relative human preference between each pair of generations, the calibration of $r$ will remain unchanged, making $\mathcal{C}_{r,{\pi_{\rm ref}}}$ more robust to the learning of $r$.
@@
The lemma provides us a unified view of the output from a language model through the space of calibrated reward.
4.2 Reward transformation for calibrated inference-time procedure
@@
An inference-time procedure $\mathcal{T}$ is called a calibrated procedure if there exists a mapping function $g_{\mathcal{T}}:[0,1]\rightarrow\mathbb{R}$ such that for any $\pi$, $r$, and ${\bm{x}},{\bm{y}}$, we have
$\mathcal{T}_{\pi}({\bm{y}}\mid{\bm{x}})\propto\pi({\bm{y}}\mid{\bm{x}})\cdot g_{\mathcal{T}}(\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}})).$
Our next result shows that for calibrated inference-time procedures and continuous language models, the aligned policy obtained from InfAlign-CTRL with any $\Phi$ has a win rate and KL divergence independent of $\pi_{\rm ref}$} and $r$.
@@
If $\mathcal{T}$ is a calibrated inference-time procedure, for any continuous language model $\pi$, $\beta>0$ and transformation function $\Phi$, we have that both $W^{\mathcal{T}}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})$ and $D_{\text{KL}}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\|{\pi_{\rm ref}})$ are independent of $r$ and $\pi_{\rm ref}$}.
The above theorem allows us to evaluate a transformation $\Phi$ by focusing on simple continuous language models that are easy to compute and simulate. In the next section, we focus on Best-of-$N$ and Worst-of-$N$, as examples to demonstrate how the theorem enables us to efficiently evaluate different transformations in practical scenarios.
4.3 Finding reward transformations for BoN and WoN
Best-of-NN inference-time procedure (BoN). During inference, $N$ i.i.d. responses from a policy $\pi$ are generated. The final output is the one with the highest reward, i.e.,
${\bm{y}}_{\texttt{BoN}}=\arg\max_{{\bm{y}}\in\{{\bm{y}}_{1},\ldots{\bm{y}}_{N}\}}r({\bm{x}},{\bm{y}}).\vskip-7.22743pt$
Worst-of-NN inference-time procedure (WoN). $N$ i.i.d. responses from a policy $\pi$ are generated, and the final output is the one with the lowest reward. i.e.,